OpenStudy (anonymous):

Use the Binomial Theorem to expand and simplify the expression. (2x - y)4

OpenStudy (dumbcow):

(2x)^4y^0 - 4(2x)^3y^1+6(2x)^2y^2 - 4(2x)^1y^3 + (2x)^0y^4

OpenStudy (anonymous):

A. 16x4 - 32x3y + 24x2y2 - 8xy3 + y4 B. 16x4 - 8x3y + 4x2y2 - 2xy3 + y4 C. 16x4 + 4x3y + 6x2y2 4xy3 + y4 D. 8x4 - 16x3y + 12x2y2 - 4xy3 + y4 E. 8x3 - 12x2y + 6xy2 - y3

OpenStudy (anonymous):

i got as far as you did but had a few things wrong.... how do I simplify it more?

OpenStudy (dumbcow):

just do the exponents and multiply ex: (2x)^2 = 2^2x^2 = 4x^2

OpenStudy (anonymous):

oh ok well that was easier than I was making it.... thanks!

OpenStudy (dumbcow):

no problem ;)